### v1

# Generating, Fast and Slow: Scalable Parallel Video Generation with Video Interface Networks

Abstract
Diffusion Transformers (DiTs) can generate short photorealistic videos, yet directly training and sampling longer videos with full attention across the video remains computationally challenging. Alternative methods break long videos down into sequential generation of short video segments, requiring multiple sampling chain iterations and specialized consistency modules. To overcome these challenges, we introduce a new paradigm called Video Interface Networks (VINs), which augment DiTs with an abstraction module to enable parallel inference of video chunks. At each diffusion step, VINs encode global semantics from the noisy input of local chunks and the encoded representations, in turn, guide DiTs in denoising chunks in parallel. The coupling of VIN and DiT is learned end-to-end on the denoising objective. Further, the VIN architecture maintains fixed-size encoding tokens that encode the input via a single cross-attention step. Disentangling the encoding tokens from the input thus enables VIN to scale to long videos and learn essential semantics. Experiments on VBench demonstrate that VINs surpass existing chunk-based methods in preserving background consistency and subject coherence. We then show via an optical flow analysis that our approach attains state-of-the-art motion smoothness while using 25-40% fewer FLOPs than full generation. Finally, human raters favorably assessed the overall video quality and temporal consistency of our method in a user study. All videos shown in this manuscript can be found here.
Figure 1: (a) Concept: Video generation has traditionally proceeded with the computationally expensive full generation or the autoregressive alternative that requires repeated inference through a sampling chain. Instead, we formulate a parallel inference paradigm: Video Interface Networks (VINs). (b) VINs enable the generation of long dynamic realistic videos that (c) often stagnate in the full attention setting or lose temporal coherence in the autoregressive setting. (d) Our method uses fewer FLOPs and achieves greater temporal consistency as captured by the Motion Aware Warped Error [17].
1 Introduction
*Work partially done during an internship at Adobe Corresponding Author {yuliu@adobe.com}
Diffusion Transformers (DiTs) [34] have heralded a revolution in generative modeling and have been scaled up to generate images [5], videos [3, 9, 44, 54], and 3D assets [31] at unprecedented fidelity. Training DiTs on longer input contexts yields remarkable generalization, albeit imposing quadratic computational and memory bottlenecks. The implications are particularly apparent in the case of high-dimensional modalities like video, where training and inference with DiTs on long videos ($>$ 64 frames) are slow and expensive [9] (see Fig 1(a), Full Denoising). Furthermore, base models trained on short videos often show motion stagnation and repetition when extended to synthesize longer videos [2, 17]. Practitioners circumvent these constraints by generating video frames in temporal chunks using the previously generated frames as a context window for the next chunk [19, 1] proceeding in a left-to-right sequence (Fig. 1(a), Autoregressive Denoising)
Generating videos sequentially is challenging because, at any given instance, the model only has a finite context. As a result, it may be susceptible to catastrophic forgetting and fail to preserve object coherence and temporal consistency in output generation. Past works [17, 6, 50] have improved the robustness of autoregression by augmenting it with long-term memory modules and image embeddings to enforce semantic and content continuity. Most techniques rely on handpicked anchors and it is not evident that a static prior suffices for generating dynamic videos. Is sequential generation the only viable approach to generating video chunks? If not, can we devise an alternative paradigm?
We draw inspiration from accounts of human cognition [42, 12, 22] that suggest that human planning adaptively switches between dual modes of ‘fast’ and ‘slow’ thinking: (1) fast effortless System 1 processing that quickly abstracts out the essence without focusing on the specifics, and (2) slow effortful System 2 processing that uses the intuited semantics to deliberate over finer details. Consider how painters approach their work – they begin by conceptualizing the overall composition through a rough sketch (System 1) before filling in colors for different objects (System 2). This natural parallelization of abstraction and generation (Fig 1(a), Parallel Denoising) remains elusive to DiTs, which solely rely on System 2-like sequential generation of low-level patches, lacking System 1 abstractions.
In this work, we introduce a video-generation paradigm that explicitly augments DiTs with a System 1-like module we call Video Interface Networks (VINs). The central insight behind VINs is simple – at each diffusion timestep, VINs encode meaningful semantics from noisy input videos into a finite set of global tokens that enable scaling to longer videos. Then, the coarse-grained global tokens guide DiTs in denoising fine-grained local tokens of disparate video chunks. The interplay of System 1 via VINs and System 2 via DiTs enables efficient parallel generation of long photorealistic videos (see Fig. 1(b)). Unlike full attention settings that diminish motion and sequential generation that obscures temporal coherence, VINs preserve both the crucial elements (Fig. 1(c)). Prior works on parallelizing video generation models use pre-specified templates, like noise rescheduling [37] or band-pass filtering [29], to enforce consistency. Instead, we train our proposed framework end-to-end, which naturally yields consistency via emergent global tokens. Our core contributions are as follows.
  • We develop a hierarchical learning paradigm that combines abstraction with generation to enable parallel generation of temporally consistent video chunks. We propose the VIN architecture (Section 3.2) to encode essential semantics of long videos. We then formulate a training objective (Section 3.3) to tightly couple the abstraction abilities of VINs with the generative abilities of DiTs and scale the joint architecture to long videos.
  • We evaluate our framework against other state-of-the-art autoregressive and parallel chunk-generation techniques [1, 17, 37, 29]. Experiments on VBench [20] (Section 4.3) demonstrate that VINs best preserve subject/background consistency across chunks and reduce the performance gap to computationally-expensive full generation. Optical flow analysis (Section 4.4) shows VINs enable pixel-level motion smoothness. VINs not only attain state-of-the-art Motion Aware Warp Error (MAWE) scores but also exhibit a superior consistency-dynamic degree tradeoff compared to the full generation at different video lengths.
  • We conduct a user study (Section 4.6) that bridges the subjectivity gap between our analysis and human ratings, demonstrating the human preference for VIN-generated videos across overall appearance and temporal consistency evaluations.
2 Related Work
Figure 2: Overview of our method. At each diffusion timestep, the noisy sample is encoded by the VIN to generate global tokens (left, Section 3.2). The VIN is trained end-to-end with the DiT by learning to denoise temporal chunks of the noisy input using the global tokens (middle, Section 3.3). At inference time, The VIN-DiT ensemble can run parallel inference on long videos (right, Section 3.4).
Diffusion Architectures for Videos. Latent Diffusion models [43, 18, 40, 19] have become the predominant approach for generating videos from text or image prompts. Early works [2, 1, 47, 14, 16, 8] focused on temporally extending pre-trained models by interleaving temporal attention with spatial attention. Subsequently, at the outset of DiTs [34], these modeling choices were simplified to denoise a sequence of video patches using Transformers [15, 30, 3, 9, 44] and yielded an unprecedented level of generalization. Training of longer sequence Transformer models can be enabled by parallelization techniques [39, 53, 41, 26], but they require massive computational resources and hardware-aware optimization. In contrast, our VIN amortizes the scaling cost into encoding global tokens.
Autoregressive Video Generation. Base models [14, 1] trained on short videos noticeably deteriorate when extended to longer videos. An important line of work [19, 1, 32, 17, 6, 50] has therefore focused on autoregressively generating temporal chunks of long videos. Sequential generation is slow, requiring each video chunk to be denoised over multiple sampling steps. Moreover, to mitigate forgetting, these methods need conditioning for global consistency, such as long-term memory modules [17], image embeddings [50, 48, 52, 6] or specialized priming [13, 51]. Pre-specified anchors increase controllability but limit the content and dynamics, are unscalable, and need specialized encoding architectures. In contrast, VINs maintain dynamic global priors and can be scaled to entire video.
Parallel Video Generation. Orthogonal to autoregressive generation, recent works [28, 37, 46, 29] have proposed generating video chunks in parallel. The central challenge lies in designing global priors that enforce inter-chunk consistency. VideoDrafter [28] generates distinct image priors for each chunk, causing temporal inconsistency. Training-free methods like FreeNoise [37] and FreeLong [29] introduce consistency priors through a noisy input and frequency filtering, respectively. Still, such priors primarily operate on surface-level visual features and fail to capture higher-level semantic abstractions. Our VINs explicitly learn deep features for global semantics at every diffusion timestep.
Learning to Abstract for Generation. Multiple approaches [21, 36, 35] integrate learning semantic encoders with the denoising objective to guide the diffusion process. Our approach is inspired by Recurrent Interface Networks (RINs) [21] that offload pixel-level denoising to learned semantic tokens. VINs combine the scalability of RIN-like architectures with DiTs. Another approach [35] trains an encoder concurrently with the denoiser to produce semantic codes for efficient diffusion but is limited to images.
3 Methodology
Our overall goal (see Fig. 2) is to (1) encode long videos into a finite set of global tokens, (2) train DiTs to generate short video chunks using the encoded tokens, and (3) combine the architectures to run parallel inference.
3.1 Preliminaries
Diffusion Models learn to generate data by reversing a gradual noising process over $T$ timesteps. The central learning objective involves learning a neural network $\epsilon_{\theta}(x_{t},t)$ to predict noise $\epsilon_{t}$ from noisy input $x_{t}$. The model $\epsilon_{\theta}(x_{t},t)$ is trained via the following regression loss:
$\mathcal{L}_{\theta}=\mathbb{E}_{t\sim\mathcal{U}(0,T),\epsilon_{t}\sim\mathcal{N}(0,1)}\left[\left\|\epsilon_{\theta}\left(x_{t},t\right)-\epsilon_{t}\right\|^{2}\right].$ (1)
Diffusion Transformers model noisy input videos as a sequence of space-time patches. For a noisy video $x\in\mathbb{R}^{H\times W\times F\times C}$, let us denote $H$ as height, $W$ as width, $F$ as frames, and $C$ as input channels. Disjoint voxels of size $(p_{1},p_{2},p_{3})$ are projected onto $N$ tokens $X^{1:N}\in\mathbb{R}^{H/p_{1}\times W/p_{2}\times F/p_{3}\times d}$ by a feedforward network $G_{\theta}(.)$. At denoising time-step $t$, the DiT passes the sequence of tokens with the text prompt embedding $T_{emb}$ through a stack of $L$ standard transformer blocks $\mathcal{T}_{\theta}^{1:L}$:
$X^{1:N,l}_{t}=\mathcal{T}_{\theta}^{l}\left(\left[X^{1:N,l-1}_{t},T_{emb}\right],t\right)\forall l\in{1,\cdots,L}.$ (2)
The final token representations $X^{1:N,L}$ are projected back to input shape by feedforward network $H_{\theta}(.)$ to obtain noise prediction $\hat{\epsilon}_{t}$. DiTs denoise inputs $x_{t}$ via quadratic self-attention on input representations, but scaling to longer videos is computationally prohibitive. We circumvent this cost using VINs, whose central computational processing is decoupled from the input.
Figure 3: VIN Architecture. The VIN encoder samples key-frames from the input video every $T_{s}$ seconds. Then, the global tokens encode the sub-sampled video via cross-attention. A stack of $M$ VIN Processor blocks refines the global tokens using self-attention.
Figure 4: Training and inference configuration. (a) During training, the VIN and DIT are trained end-to-end to denoise video chunks. (b) At inference time, the architectures are frozen to run parallel denoising with token fusion. SG denotes the stop gradient operation.
3.2 Video Interface Networks
VINs condense videos with arbitrary lengths into essential semantics (see Fig. 3). VINs encode the inputs into a set of global tokens via the VIN encoder. Then, VIN models these tokens via consecutive blocks of the VIN processor. We describe each of the components in detail next.
Global Tokens are at the heart of the VIN backbone. They are instantiated as fixed-sized embeddings $Z_{init}\in\mathbb{R}^{N_{global}\times d}$, independent of the input.
VIN Encoder encodes the input into the global tokens. First, we sub-sample the input video $X_{t}^{1:N}$ at fixed time intervals $T_{s}$ to obtain the key-frames $X_{t}^{1:N,T_{s}}$. Then, the encoder uses the multi-head attention (MHA) mechanism to read the sub-sampled video into $Z_{init}^{1:N_{global}}$. More specifically, we derive the query vectors $Q$ from $Z_{init}^{1:N_{global}}$ and key-values vectors from $X_{t}^{1:N,T_{s}}$. The query-key-value interaction enables the selective encoding of the input in the global tokens to obtain input-aware instantiation $Z_{t}$:
$Z_{t}$ $=MHA\left(Q=Z_{init};\;K,V=X_{t}^{1:N,T_{s}}\right)$ (3)
VIN Processor iteratively refines the instantiated tokens via self-attention. In addition, latents are also conditioned on prompt $T_{emb}$:
$Z_{t}$ $=MHA\left(Q,K,V=\left[Z_{t},T_{emb}\right]\right)$ (4)
We stack $M$ blocks together such that the bulk of the VIN computation is allocated to refining the global tokens.
3.3 Learning Objective
DiTs and VINs have orthogonal strengths – DiTs excel at modeling fine-grained details in short video segments, while VINs can filter redundancy to learn higher-level features from longer videos. Our strategy is to formulate a learning objective that couples the complementary abilities of DiTs and VINs to generate long videos. To this end, we first integrate the global tokens in the noise estimate of denoising step $t$:
[…]
Next, we relax and factorize the noise distribution across temporal chunks. Let the video with $F$ frames have $N$ tokens. Correspondingly, each chunk with $F_{chunk}$ frames contains $N_{S}$ tokens. The resulting factorization is given by:
$P_{\theta}(\epsilon_{t}|X_{t},t,Z_{t})=\\ \prod_{i=0}^{\lfloor{N/N_{s}}\rfloor}P_{\theta}(\epsilon_{t}^{iN_{s}:(i+1)N_{s}}|X_{t}^{iN_{s}:(i+1)N_{s}},t,Z_{t})$ (6)
We parametrize $P_{\theta}(.)$ as a DiT network $\epsilon_{\theta}(.)$ and $P_{\alpha}(.)$ as a point estimate from the VIN network $\delta(f_{\alpha}(.))$. The learning objective resulting from the factorized formulation is as follows:
$Z_{t}=f_{\alpha}(X_{t},Z_{init}^{1:N_{global}})$ (7)
$\mathcal{L}_{\alpha,\theta}=\mathbb{E}_{\epsilon_{t}\sim\mathcal{N}(0,1)}[\sum_{i=0}^{\lfloor{N/N_{s}}\rfloor}\left\|\epsilon_{\theta}\left([X_{t}^{i},\;Z_{t}],t\right)-\epsilon_{t}^{i}\right\|^{2}\ ]$ (8)
where $X_{t}^{i}\equiv X_{t}^{iN_{s}:(i+1)N_{s}}$ and $\epsilon_{t}^{i}\equiv\epsilon_{t}^{iN_{s}:(i+1)N_{s}}$ denote the $i^{th}$ chunk of the input and noise estimate, respectively. Eq. (8) enables the DiT to simultaneously train on different temporal chunks in parallel. Moreover, end-to-end training enables the architectures to bootstrap off each other where (1) the DiT learns to ground denoising in encoded tokens and (2) the VIN learns to refine global tokens to enable downstream denoising. In practice, neighboring chunks have useful local contexts for denoising. During training, for each chunk $i$, we also append $N_{local}$ tokens from the last $F_{local}$ frames of the previous chunk $i-1$ to the DiT.
$\hat{\epsilon}_{t}^{i}=\epsilon_{\theta}\left([X_{t}^{i},\;Z_{t}\;,SG\left(X_{t}^{i-1,N_{local}}\right)],t\right)\\ \forall i\in 0,\cdots,\lfloor{N/N_{s}}\rfloor$ (9)
Here, $SG(.)$ stands for the stop gradient operation. We observed that sharing gradients between chunks worsens model output fidelity. Hence, we use $SG(.)$ to prevent inter-chunk gradient flow. Fig. 4(a) visualizes the training step at diffusion timestep $t$.
3.4 Inference Method
Inference at time-step $t$ proceeds by (1) encoding the current estimate $x_{t}$ into the global tokens $Z_{t}$ via VINs and (2) using $Z_{t}$ to denoise chunks parallelly via the DiT to obtain the next sample $x_{t-1}$. As outlined in Eq. (9), the adjacent chunks contain overlapping local context and the shared context may diverge across chunks after individual denoising. We mitigate the divergence by fusing the overlapping tokens (see Fig. 5) of noise prediction $\hat{\epsilon}_{t}$. For each position $k$ in the overlap region between chunks $\hat{\epsilon}_{t}^{i}$ and $\hat{\epsilon}_{t}^{i+1}$, we fuse the overlapping tokens through the following weighted average:
$\hat{\epsilon}_{t}^{\text{fused}}[k]=\frac{\left(\mathcal{F}_{local}-\mathcal{W}(k)\right)\hat{\epsilon}_{t}^{i}[k]+\mathcal{W}(k)\hat{\epsilon}_{t}^{i+1}[k]}{\mathcal{F}_{local}}$ (10)
Here, $\mathcal{W}(k)\in\{1,\cdots,\mathcal{F}_{local}\}$ denotes the relative temporal position of the token in the overlapping region. We also investigate fusion schedules in the sampling chain as follows: (a) early fusion: token fusion applied in initial stages $t>t_{\alpha}$, (b) mid fusion: fusion constrained to specific sampling interval $t_{\alpha}<t<t_{\beta}$, and (c) late fusion: fusion implemented exclusively in latter steps $t<t_{\alpha}$. Fig. 4(b) visualizes the inference at diffusion timestep $t$ along with token fusion. Appendix 11 details the joint VIN and DiT training and inference algorithms.
Figure 5: Inference Time Token Fusion. Overlapping tokens in neighboring chunks are combined by a weighted average of their relative temporal position in the chunk.
4 Experiments
We assess the capability of VINs to generate video chunks with temporal consistency and semantic coherence, benchmarking it against state-of-the-art sequential and parallel generation approaches.
4.1 Experimental Setup
Base Model: Our method fine-tunes a pre-trained latent video DiT based on a modified Open-Sora variant [54]. The base model encodes 16 video frames into five latent frames using a 3D Variational Auto-Encoder [25]. The final videos are generated at $192\times 320$ resolution and 16 FPS.
VIN: We augment the base model with the proposed VIN backbone. The architecture uses one encoder and $M=4$ processor blocks, modeling $N_{global}=512$ tokens with 32 attention heads and hidden dimension $d=4096$. The VIN encoder samples input frames every $T_{s}=1.0s$.
Training Dataset: Our models are trained on 840000 captioned videos. These videos are licensed and have been filtered to remove low-quality content. The training set consists of a mix of $\{64,128,256\}$ frames or $\{20,40,80\}$ latent frame videos.
Training Details: The VIN and DiT ensemble are trained end-to-end (Section 3.3) to parallelly generate $F_{chunk}=20$ latent frame video chunks. Each chunk receives $F_{local}=8$ local latent frames and $512$ global tokens.
Inference Details: At inference time, we run reverse diffusion for $50$ timesteps using an expanded $F_{local}=12$. We find and later demonstrate through ablations that early token fusion is the most effective and set fusion cutoff ($t_{\alpha}=20$).
Test Set: We evaluate models on text-to-video generation from a diverse swath of prompts. Our prompt set comprises 150 prompts (Appendix 13) selected by randomly sampling 25 test prompts across six VBench [20] dimensions viz. [subject consistency, background consistency, temporal flickering, temporal style, motion smoothness, overall consistency]. We generate videos at four different lengths, 64, 128, 256, and 512 frames, and sample three videos per prompt for each length to evaluate model performance.
Figure 6: VBench evaluation results. We perform comparisons based on (a) subject consistency, (b) background consistency, (c) imaging quality and (d) dynamic degree. VINs demonstrate superior consistency compared to other chunk-based methods and preserve dynamic degree as opposed to full attention. (d) Visual examples compare VINs against the autoregressive and spectral blending methods on common prompts. VIN samples possess greater semantic consistency, subject identity preservation, and stable backgrounds across frames.
4.2 Comparison Methods
Next, we compare VINs to methods across a broad class of video chunk generation paradigms.
Full Attention (Full): The pre-trained model generates the entire video simultaneously by attention across all tokens.
Autoregressive (AutoReg.): Following prior work [19, 1], we use the replacement with noise method to fine-tune our base model to generate chunks sequentially. We generate $10$ latent frames for every step and use $15$ context frames.
StreamingT2V (ST2V): This state-of-the-art sequential generation technique [17] trains memory modules to facilitate temporal coherency. We compare VINs against their text-to-video implementation  [^1] StreamingModelscope Github Repository.
FreeNoise: This parallel inference method [37] uses noise rescheduling and temporal fusion to generate video chunks in parallel. We apply it to our base model with a 20-frame chunk size and 12-frame fusion window.
Spectral Blending (Spec. Blend): This state-of-the-art parallel inference technique [29] uses band-pass filtering to blend video chunks together. We implemented it on our base model to blend 20-frame chunks together.
4.3 VBench Evaluation
Disparate generations of video chunks must preserve temporal consistency, maintaining long-range dependencies across chunks. To this end, we use the VBench Long  [^2] VBench Long Github Repository [20] benchmark to evaluate the long-range consistency and quality of generated videos via four key metrics:
  1. 1.
    Subject Consistency: Temporal object consistency using DINOv2 features [33] (‘fast arm’ of VBench Long).
  2. 2.
    Background Consistency: Temporal background consistency using CLIP [38] features (‘fast arm’).
  3. 3.
    Imaging Quality: Image distortion (over-exposure, blur) using MUSIQ [23].
  4. 4.
    Dynamic Degree: Motion magnitude using RAFT optical flow magnitudes [45].
Fig. 6(a), (b), (c) and (d) show the evaluation results across the four dimensions. The major takeaways are as follows:
VINs improve long-range temporal consistency. VINs improve subject and background consistency across video chunks, as measured by DINO and CLIP features (Figs. 6(a), (b)), reducing catastrophic forgetting compared to ST2V and an autoregressive baseline. Moreover, VIN meaningfully closes the gap with full generation.
Full Attention not only stagnates at longer lengths but paradoxically performs worse despite its quadratic complexity, as evidenced by the deterioration of the Full method in the dynamic degree plot (Figs. 6 (d)). In contrast, VINs maintains high dynamic degree and temporal consistency throughout long videos while remaining efficient.
VINs trade off temporal consistency for subject-aware softening. We observed that while VINs maintain subject fidelity, they exhibit a marginal difference in image quality, as reflected in the scores of Fig. 6(c). However, the quality remains competitive with other techniques, also validated by the user study in Section 4.6.
Visual examples. Fig. 6(e) illustrates our observations via four diverse prompts. We notice that VIN-generated videos have greater subject and background stability and preserve the relative positions of the objects in the scene. See Appendix 7 for more evaluations and Appendix 8 for qualitative visualizations.
Figure 7: Optical Flow Analysis: (a) VINs demonstrate more consistent and natural-looking movement. They outperform the MAWE scores of the state of the art by a significant margin. (b) Per-frame optical flow analysis shows that VIN-generated videos simultaneously have low warp error and considerable optical flow, implying smooth and dynamic videos. (c) In addition, VINs also reduce abrupt scene changes. (d) Visualizing $y$-$t$ cross-section of generated videos reveals motion smoothness endowed by VIN.
4.4 Optical Flow Analysis
While VINs demonstrate consistency as measured by deep neural network features, such a consistency should also reliably transfer to pixel-level generation. To this end, this section analyzes the optical flow across generated frames with the RAFT model [45].
Metrics. We adopt the MAWE metric [17], which, for a video $V$, is defined as $W(V)/(c\times OFS(V))$, where $W(V)$ measures the average squared L2 pixel distance between a frame and its corresponding warped frame, excluding occluded regions, $OFS(V)$ computes the mean magnitude of optical flow vectors between consecutive frames, and $c=9.5$ is the calibration constant. A low MAWE score suggests that the generated video exhibits temporal coherency, with high optical magnitudes. In addition, we also use PySceneDetect [4] to detect abrupt scene changes and report Scene-Cut Rates per video. Fig. 7 reports the results of the analysis. We highlight some key observations.
VINs outperform previous techniques in terms of motion consistency. Fig. 7(a) shows that VINs maintain a considerably lower MAWE across different video lengths, with scores consistently remaining below 2.0 even at 512 frames, in contrast to baseline methods. Note that other methods that use the same base model do not exhibit this improvement. Fig. 7(b) shows frame-wise MAWE, warping error, and optical flow. VINs exhibit stable optical flow (1.5-2.0) and consistently lower warping errors (4.0-5.0) compared to others, which have more erratic and bursty motion.
VINs maintain scene continuity across video chunks at longer lengths. The heatmap in Fig. 7(c) demonstrates that, compared to other chunking methods, VINs achieve the lowest scene-cut detection rates for longer video lengths of $256$ and $512$ frames, minimizing the performance gap between chunked and direct video generation. Fig. 7(d) visualizes the $y$-$t$ cross-section of $256$ frame videos generated by VINs and the Autoregressive method. VINs demonstrate smoother transitions with fewer aberrations.
Consistency across chunk transitions. We specifically analyzed consistency across chunk boundaries in Fig. 8 by measuring the warp error and MAWE around chunk transitions. In addition, we estimated the gold standard baseline by measuring transition errors from ground truth decoded latents obtained from the 3D VAE (Oracle). Our method achieves comparable warp error to full attention and significantly reduces the MAWE gap with the Oracle. Fig. 9 visualizes the smooth chunk transitions on generated videos.
Figure 8: Optical flow analysis of chunk transitions in 256 frame videos. Measured over $\Delta=16$ frames before chunk boundary. Numbers above the bars show gap from the 3D VAE Oracle.
Figure 9: Chunk Transition Visualization. Generated videos show a chunk counter on the top left. A red dot appears on the top right before chunk boundaries.
Figure 10: Comparison to full attention: (a) VIN maintains a superior consistency-dynamic degree tradeoff across different frames. (b) Efficiency analysis of the full vs. VIN model. Latency and memory are measured on NVIDIA A100 80GB GPU using the same memory-efficient transformer implementation.
Comparison to Full Attention. Video generators face a natural tradeoff between temporal consistency and motion dynamism. Our analysis in Fig. 10 (a) demonstrates that VINs achieve a superior balance in this tradeoff, particularly as frame counts increase. Naively extending DiT context with full attention diminishes motion and causes repetitions due to diluted temporal connections (see Appendix 8). Beyond motion quality, our approach addresses critical scalability limitations of full attention. As illustrated in Fig.10 (b), VINs deliver 25-40% higher efficiency (also Fig. 1) and 40-75% faster denoising speeds while requiring only marginally more memory, making them significantly more practical for longer video generation.
4.5 VIN Ablations and Interpretability
Ablations. Table 1 presents the results of ablation experiments on key VIN architectural components. The full model achieved the best overall performance, while removing global tokens significantly reduced temporal consistency. Past works [24] have posited that diffusion models form object-structure features early in the sampling process; thus, our full model benefited from early fusion compared to other fusion schedules. Expanding the local context at inference time also improves motion flow. Interestingly, increasing the number of keyframes sampled does not improve consistency, suggesting redundancy of dense temporal sampling for VINs.
VIN Interpretability. We interpret learned VIN representations by visualizing the attention distribution across input video keys (Fig. 11). The distributions revealed semantically meaningful attention patterns, with heads focusing on human forms, architecture, and objects. See Appendix 10 for attention across entire videos.
Model Configuration MAWE $\downarrow$ Scene Cuts $\downarrow$
Full Model 1.09 0.21
w/o Global Tokens 1.69 0.33
w/o fusion 1.13 1.00
Mid fusion $(t_{\alpha},t_{\beta}=35,15)$ 1.11 0.33
Late fusion $(t_{\alpha}=20)$ 1.22 0.74
Local Frames - 8 \10 1.51 \1.17 0.24 \0.22
Keyframe $T_{s}$ 0.5s \0.2s 1.14\1.21 0.34 \0.29
Table 1: Effect of model ablations on motion consistency.
Figure 11: VIN Interpretatibility. VIN attention heads focus on semantically meaningful objects in the scene.
4.6 User Study
We set up a user study (design outlined in Appendix 12) where a cohort of humans was asked to compare videos generated by VINs against other methods on (1) overall appearance and (2) temporal consistency. Fig. 12 reports the percentage results with which VIN is deemed better, comparable, or worse. In contrast to the imaging quality results of Section 4.3, humans rated VIN-generated videos to be comparable on average or even better (loss rate $<$ 30 %). Moreover, the judged temporal consistency largely correlates with the VBench results, where VIN preference matches full generation, substantially improves upon parallel methods, and outperforms sequential generation.
Figure 12: User study results.
4.7 Discussion
Dynamic universal representations enable long-range video consistency. Unlike context-limited autoregressive methods [19, 1] or static priors [17], VINs use global tokens recomputed each step, degrading gracefully with video length. Deep hierarchical priors improve motion smoothness. Compared to surface-level priors (FreeNoise [37], Spectral Blending [29]) and DiT, VINs achieve lower MAWE scores, suggesting deep semantic representations enable fluid long-video motion. Limitations: VINs only learn through generation. Future work could explore concurrent learning on downstream tasks for richer representations and depth/3D input beyond raw patches.
5 Conclusion
We proposed a paradigm to parallelize inference on videos by using global tokens from long videos to denoise shorter video chunks. Our experiments show that this hierarchy improves subject and background consistency across chunks, endows fluid motion, and surpasses other state-of-the-art chunking techniques. Compared to full attention, VIN improves the consistency-motion dynamism tradeoff besides being scalable. The generated videos are also judged favorably by human raters.